OpenStudy (pawanyadav):

Lim{(x+x^-1)e^(1÷x)-x=? Where x tending to infinity.

geerky42 (geerky42):

\[\lim_{x\to\infty} (x+x^{-1})e^{1/x}-x\]?

OpenStudy (pawanyadav):

Ya this is the question. what's the solution?

OpenStudy (welshfella):

Hint : as x approaches infinity 1/x approaches zero so e^1/x will approach 1.

geerky42 (geerky42):

\[\lim_{x\to\infty} (x+x^{-1})e^{1/x}-x \\~\\= \lim_{x\to\infty} xe^{1/x}+x^{-1}e^{1/x}-x\\~\\ = \lim_{x\to\infty} x(e^{1/x}-1)+x^{-1}e^{1/x}\] Here, \(x^{-1}e^{1/x}\) approaches to 0. So we are left with \[\lim_{x\to\infty} x(e^{1/x}-1)\\~\\=\lim_{x\to\infty} \dfrac{e^{1/x}-1}{\dfrac{1}{x}}\]From here, use L'Hopital's Rule.

geerky42 (geerky42):

@Pawanyadav Does that help?

OpenStudy (anonymous):

I don't think so because it's not the solution :P

OpenStudy (pawanyadav):

Yes absolutly

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

geerky42 (geerky42):

Correct.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

geerky42 (geerky42):



OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

OpenStudy (pawanyadav):

The answer After solving is 1.

geerky42 (geerky42):

@Zarkon I think something is buggy.

geerky42 (geerky42):

I see so many replies "The answer After solving is 1."